Show that f(x)=2/x+1 is indected


1
Expert's answer
2021-02-24T06:59:03-0500

"f(x)=\frac{2}{x+1}"


"y'=-\frac{2}{(x+1)^2}"


"y''=\frac{4}{(x+1)^3}"


For "x\isin (-\infin,-1)" : "y''<0\implies f(x)" is concave downward.

For "x\isin (-1,\infin)" : "y''>0\implies f(x)" is concave upward.